40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4882     Entry time: Sat Jun 25 00:00:28 2011
Author: Sonali 
Type: Update 
Category: Green Locking 
Subject: Fibre Coupling. 

 What I did today.

1. I tried to align the IR input beam by aligning the two mirrors, to couple input light into the fibre.

2.I was unsuccessful for a long time even though I tried a lot of tricks.

3. I also tried to use the optical fault locator to superpose the IR beam spot onto the beam spot of the other laser to facilitate effective coupling.

4.But the crucial point was to superpose the input beam path in the perfect direction of the output beam path and not just the beam spot.(the input cone and the output cone are perfectly aligned).

5.After one whole day of trial and thought, I managed to couple light into the fibre, and saw the output beam spot on the screen-camera-monitor set-up which we had arranged. Eurekka !!;)

6.I then used a power meter to measure the input beam power and the output beam power.

7.It was a disappointing 2% . I had read in project reports of many students of a 20% success.

8.After a lot of subtle tweaking of the mirrors using the knobs, I managed to increase the percentage of output beam to 12%.

9. This is a workable level.

10.A day of lot of new learning! Pictures of the setup are attached.:)

 

Attachment 1: Fibre_coupling_successful_24_june.jpg  2.974 MB  | Hide | Hide all
Fibre_coupling_successful_24_june.jpg
Attachment 2: Beam_output_on_screen.jpg  1.777 MB  | Hide | Hide all
Beam_output_on_screen.jpg
ELOG V3.1.3-